# Singular Value and Frame Decomposition-based Reconstruction for Atmospheric Tomography

### v1

@@ 3.1 Computational Aspects
$\phi_{\alpha}^{\delta}:=\sum_{j,k\in\mathbb{Z}}\left(\sum_{n=1}^{r^{(s)}_{jk}}g_{\alpha}\left((\sigma^{(s)}_{jk,n}\right)\left((u^{(s)}_{jk,n})^{H}\varphi^{\delta}_{jk}\right)v^{(s)}_{jk,n}\right)_{\ell}w^{(s)}_{jk,\ell}(x,y)\,.$ (3.10)
@@
Note that in Algorithm 1 we have implicitly assumed that both the atmospheric turbulence $\phi$ and the wavefronts $\varphi$ are periodic functions defined on the square domain $\Omega_{T}$. In practice, wavefronts are typically only given on a subset $\Omega_{A}\subset\Omega_{T}$ corresponding to the telescope aperture, and thus need to be extended outside $\Omega_{A}$. The effect of this extension was studied in [Gerth_2014], where it was found to produce only very minor errors in the reconstructions along the boundaries of the domains $c_{\ell}\Omega_{T}$.
@@ 4 An Explicit Frame Decomposition
The definition (4.7) of the frame inverse $\mathcal{A}$ involves the dual frame functions $\widetilde{w}_{jk,\ell g}$, which had to be computed numerically in [Hubmer_Ramlau_2020, Weissinger2021]. However, we now show that these functions $\widetilde{w}_{jk,\ell g}$ have a simple analytic expression, which in turn leads to an explicit representation of $\mathcal{A}$ that does not involve infinite sums. For this, let
@@
with $I_{\ell g}$ as in (4.5) denote the overlay functions of the $\ell$-th atmospheric layer.
@@
=(2.5)${=}\sum_{g=1}^{G}\sum_{j,k\in\mathbb{Z}}\left\langle\,f,w_{jk,l}I_{\ell g}\,\right\rangle_{L^{2}(\Omega_{\ell},\gamma_{\ell})}w_{jk,l}(x,y)I_{\ell g}(x,y)$ )
$=\sum_{g=1}^{G}I_{\ell g}(x,y)\sum_{j,k\in\mathbb{Z}}\left\langle\,fI_{\ell g},w_{jk}\,\right\rangle_{L^{2}(\Omega_{\ell},\gamma_{\ell})}w_{jk,l}(x,y)$
@@
$=(2T)\sum_{j,k\in\mathbb{Z}}\sum_{l=1}^{L}\frac{\sqrt{\gamma_{\ell}}}{c_{\ell,g}}w_{jk}\left(\frac{\alpha_{g}h_{\ell}}{c_{\ell,g}}\right)\left\langle\,\phi_{\ell},w_{jk,\ell g}\,\right\rangle_{L_{2}(\Omega_{\ell},\gamma_{\ell})}\left\langle\,w_{jk},\varphi_{g}\,\right\rangle_{L_{2}(\Omega_{A})}$
$=\sum_{l=1}^{L}\left\langle\,\phi_{\ell},(2T)\sum_{j,k\in\mathbb{Z}}\frac{\sqrt{\gamma_{\ell}}}{c_{\ell,g}}w_{jk}\left(-\frac{\alpha_{g}h_{\ell}}{c_{\ell,g}}\right)\left\langle\,\varphi_{g},w_{jk}\,\right\rangle_{L_{2}(\Omega_{A})}w_{jk,\ell g}\,\right\rangle_{L_{2}(\Omega_{\ell},\gamma_{\ell})}\,,$
@@ 5.1 Test Configuration and Simulation Environment
The AO system configuration is shown in Table 5.2. We simulate a telescope which gathers light through a primary mirror with $42$ m diameter and a $28\%$ central obstruction. The ELT optical design consists of three mirrors denoted by M1, M2 and M3 on-axis with two DMs (M4, M5) for performing the AO. For MORFEO, two additional DMs (DM1, DM2) inside the instrument are used for wavefront compensation. Note that we assume the Fried geometry with equidistant actuator spacing for all DMs. Details on the configuration are listed in Table 5.2. The turbulence is simulated according to median seeing conditions with a Fried parameter of $0.129$ m. We use the ESO standard 9-layer atmosphere and reconstruct $3$ layers at the altitudes of the DMs, see Table 5.4. In this way, we obtain the DM shapes directly from the reconstructed layers and do not have to perform a projection step.
@@
We test our algorithms on two different configurations of natural guide stars (NGS) and laser guide stars (LGS) for measuring the wavefront aberrations, see Figure 5.1. In the first test case (left Figure) we simulate $6$ bright NGS positioned in a circle of $1$ arcmin diameter. This setting is easier to handle, as the modelling of LGS is more complex due to their finite height. Moreover, wavefronts obtained by LGS face the problem of tip-tilt indetermination [Roddier1999], i.e., the planar component of the measured wave is unreliable. In practice, it is not realistic to find $6$ bright NGS in the surrounding of the object of interest. Thus, we consider as second test configuration (right Figure) the standard MORFEO setting with $6$ LGS positioned in a circle of $1$ arcmin diameter. We account for the tip-tilt effect via removing the planar component of LGS wavefronts using $3$ faint NGS located in a circle of $160$ arcsec diameter [Gilles_Ellerbroeck_2008]. We model the sodium layer at which the LGS beam is scattered via a Gaussian random variable with mean altitude $H=90$ km and FWHM of the sodium density profile of $11.4$ km. To each bright NGS and LGS, a Shack-Hartmann (SH) WFS with $84\times 84$ subapertures is assigned. The faint NGS for tip-tilt removal are equipped with $2\times 2$ SH WFS. A detector read-out noise (RON) of $3.0$ electrons per pixel per frame is used. For more details see Table 5.4.
@@
We examine the performance of the proposed methods against the Gradient method [Saxenhuber_Ramlau_2016, Saxenhuber_2016] and FEWHA [Yudytskiy_2014, Yudytskiy_Helin_Ramlau_2014], both being iterative reconstruction algorithms for atmospheric tomography. The Finite Element Wavelet Hybrid Algorithm (FEWHA) uses a dual-domain discretization strategy into wavelet and bilinear basis functions leading to sparse operators. A matrix-free representation of all operators involved makes FEWHA very fast and enables on-the-fly system updates whenever parameters at the telescope or in the atmosphere change [Stadler2020, Stadler2021]. The sparse system is solved using the CG method with a Jacobi preconditoner [Yudytskiy_Helin_Ramlau_2013] and an augmented Krylov subspace method [RaSt2021, Stadler2021] to reduce the number of iterations. The algorithm leads in real-time to an excellent reconstruction quality compared to standard matrix-vector multiplication approaches for the MORFEO instrument [Stadler2021, Stadler2022]. The Gradient method amounts to a steepest descent method applied to a least squares functional, employed with an accelerated step-size, developed and analyzed in [Saxenhuber_Ramlau_2016, Saxenhuber_2016], here applied with a separated tip-tilt reconstruction in the mixed case.
In the community of AO, it is common to validate the reconstruction quality using the so called Strehl ratio into certain directions. In our simulations, we use $25$ directions positioned in a $5\times 5$ grid over the FoV. The Strehl ratio is defined as the ratio between the maximum of the real energy distribution of incoming light in the image plane $I(x,y)$ over the hypothetical distribution $I_{D}(x,y)$, which stems from the assumption of diffraction-limited imaging, i.e.,
$\text{SR}:=\frac{\max_{(x,y)}I(x,y)}{\max_{x,y}I_{D}(x,y)}\,.$
@@ 5.2 Numerical Results
As a first step of performance evaluation, we consider not a whole end-to-end AO simulation, but compare only the tomographic reconstruction of the proposed methods. We simulate a $3$ layer atmosphere (top row) and perform its reconstruction with the SVTD (middle row) and the iterative FD (bottom row) methods. The results are shown in Figure 5.2. We observe that both algorithms provide a very good result at the lowest layer $\ell=1$, but fail to reconstruct details at higher altitudes, in particular around the borders. This behaviour is expected, and generally observed for atmospheric tomography, since the reconstruction quality depends on the number of overlapping regions, which decreases with height and field off-axis position [Ramlau_Stadler_2024].
@@
Since all algorithms yield promising tomographic reconstructions, we next consider full end-to-end AO simulations and evaluate the short-exposure (SE) and long-exposure (LE) Strehl ratio. The SE Strehl ratio provides a measure for the reconstruction quality at the current time step, whereas the LE Strehl ratio is an average over time. We compare the SVTD and FD approaches with FEWHA and the Gradient method, which are both known to provide very good results for the considered test configurations. For all results with the SVTD method, the regularization parameter $\alpha$ and the smoothing parameter $s$ were tuned experimentally a-priori. The results for the iterative FD and the Gradient method use $5$ iterations, whereas FEWHA uses $4$ iterations in its internal CG method. The method specific parameters for FEWHA have been tuned via simulations.

### v2

@@ 3.1 Computational Aspects
$\phi_{\alpha}^{\delta}:=\sum_{j,k\in\mathbb{Z}}\left(\sum_{n=1}^{r^{(s)}_{jk}}g_{\alpha}\left(\sigma^{(s)}_{jk,n}\right)\left((u^{(s)}_{jk,n})^{H}\varphi^{\delta}_{jk}\right)v^{(s)}_{jk,n}\right)_{\ell}w^{(s)}_{jk,\ell}(x,y)\,.$ (3.10)
@@
Note that in Algorithm 1 we have implicitly assumed that both the atmospheric turbulence $\phi$ and the wavefronts $\varphi$ are periodic functions defined on the square domain $\Omega_{T}$. In practice, wavefronts are typically only given on a subset $\Omega_{A}\subset\Omega_{T}$ corresponding to the telescope aperture, and thus need to be extended outside $\Omega_{A}$. The effect of this extension was studied in [Gerth_2014], where it was found to produce only very minor errors in the reconstructions along the boundaries of the domains $c_{\ell}\Omega_{T}$. Thus, all wavefronts used in the numerical examples of Section 5 are extended by $0$
@@ 4 An Explicit Frame Decomposition
The definition (4.7) of the frame inverse $\mathcal{A}$ involves the dual frame functions $\widetilde{w}_{jk,\ell g}$, which had to be computed numerically in [Hubmer_Ramlau_2020, Weissinger2021]. However, we now show that these functions $\widetilde{w}_{jk,\ell g}$ have a simple analytic expression, which in turn leads to an explicit representation of $\mathcal{A}$ that does not involve infinite sums. For this, we denote the overlay functions of the $\ell$-th atmospheric layer as
@@
with $I_{\ell g}$ as in (4.5).
@@
=(2.5)${=}\sum_{g=1}^{G}\sum_{j,k\in\mathbb{Z}}\left\langle\,f,w_{jk,\ell}I_{\ell g}\,\right\rangle_{L^{2}(\Omega_{\ell},\gamma_{\ell})}w_{jk,\ell}(x,y)I_{\ell g}(x,y)$ )
$=\sum_{g=1}^{G}I_{\ell g}(x,y)\sum_{j,k\in\mathbb{Z}}\left\langle\,fI_{\ell g},w_{jk}\,\right\rangle_{L^{2}(\Omega_{\ell},\gamma_{\ell})}w_{jk,\ell}(x,y)$
@@
$=(2T)\sum_{j,k\in\mathbb{Z}}\sum_{\ell=1}^{L}\frac{\sqrt{\gamma_{\ell}}}{c_{\ell,g}}w_{jk}\left(\frac{\alpha_{g}h_{\ell}}{c_{\ell,g}}\right)\left\langle\,\phi_{\ell},w_{jk,\ell g}\,\right\rangle_{L_{2}(\Omega_{\ell},\gamma_{\ell})}\left\langle\,w_{jk},\varphi_{g}\,\right\rangle_{L_{2}(\Omega_{A})}$
$=\sum_{\ell=1}^{L}\left\langle\,\phi_{\ell},(2T)\sum_{j,k\in\mathbb{Z}}\frac{\sqrt{\gamma_{\ell}}}{c_{\ell,g}}w_{jk}\left(-\frac{\alpha_{g}h_{\ell}}{c_{\ell,g}}\right)\left\langle\,\varphi_{g},w_{jk}\,\right\rangle_{L_{2}(\Omega_{A})}w_{jk,\ell g}\,\right\rangle_{L_{2}(\Omega_{\ell},\gamma_{\ell})}\,,$
@@ 5.1 Test Configuration and Simulation Environment
The AO system configuration is shown in Table 5.2. We simulate a telescope which gathers light through a primary mirror with $42$ m diameter and a $28\%$ central obstruction. The ELT optical design consists of three mirrors denoted by M1, M2 and M3 on-axis with two DMs (M4, M5) for performing the AO. For MORFEO, two additional DMs (DM1, DM2) inside the instrument are used for wavefront compensation. Note that we assume the Fried geometry with equidistant actuator spacing for all DMs [Fried_77]. Details on the configuration are listed in Table 5.2. The turbulence is simulated according to median seeing conditions with a Fried parameter of $0.129$ m. We use the ESO standard 9-layer atmosphere as e.g. defined in [Auzinger_2015] and reconstruct $3$ layers at the altitudes of the DMs, see Table 5.4. Figure 5.1 shows a graphical illustration of the DM and layer configuration. The layer heights are chosen such that the light emitted from a certain turbulent layer is exactly focused in the corresponding DM. A wavefront perturbation stemming from one of these layers, which are assumed to be infinitely thin, is then exactly compensated if the corresponding DM is given the correct shape. This special choice of the mirror position along the optical axis is often referred to as conjugation. In this way, we obtain the DM shapes directly from the reconstructed layers and do not have to perform a projection step.
@@
Figure 5.1: Graphical illustration of 3 DMs conjugated to 3 reconstructed layers.
We test our algorithms on two different configurations of natural guide stars (NGS) and laser guide stars (LGS) for measuring the wavefront aberrations, see Figure 5.2. In the first test case (Figure 5.2, left) we simulate $6$ bright NGS positioned in a circle of $1$ arcmin diameter. This setting is easier to handle, as the modelling of LGS is more complex due to their finite height. Moreover, wavefronts obtained by LGS face the problem of tip-tilt indetermination [Roddier1999], i.e., the planar component of the measured wave is unreliable. In practice, it is not realistic to find $6$ bright NGS in the surrounding of the object of interest. Thus, we consider as second test configuration (Figure 5.2, right) the standard MORFEO setting with $6$ LGS positioned in a circle of $1$ arcmin diameter. We account for the tip-tilt effect via removing the planar component of LGS wavefronts using $3$ faint NGS located in a circle of $160$ arcsec diameter [Gilles_Ellerbroeck_2008]. We model the sodium layer at which the LGS beam is scattered via a Gaussian random variable with mean altitude $H=90$ km and FWHM of the sodium density profile of $11.4$ km. To each bright NGS and LGS, a Shack-Hartmann (SH) WFS with $84\times 84$ subapertures is assigned. The faint NGS for tip-tilt removal are equipped with $2\times 2$ SH WFS. The noise induced by the detector read-out is simulated as $3.0$ electrons per pixel and frame. A detector read-out noise (RON) of $3.0$ electrons per pixel per frame is used. For more details see Table 5.4.
@@
We examine the performance of the proposed methods against the Gradient method [Saxenhuber_Ramlau_2016, Saxenhuber_2016] and the Finite Element Wavelet Hybrid Algorithm (FEWHA) [Yudytskiy_2014, Yudytskiy_Helin_Ramlau_2014], both being iterative reconstruction algorithms for atmospheric tomography. FEWHA uses a dual-domain discretization strategy into wavelet and bilinear basis functions leading to sparse operators. A matrix-free representation of all operators involved makes FEWHA very fast and enables on-the-fly system updates whenever parameters at the telescope or in the atmosphere change [Stadler2020, Stadler2021]. The sparse system is solved using the CG method with a Jacobi preconditoner [Yudytskiy_Helin_Ramlau_2013] and an augmented Krylov subspace method [RaSt2021, Stadler2021] to reduce the number of iterations. The algorithm leads in real-time to an excellent reconstruction quality compared to standard matrix-vector multiplication approaches for the MORFEO instrument [Stadler2021, Stadler2022]. The Gradient method amounts to a steepest descent method applied to a least squares functional, employed with an accelerated step-size, developed and analyzed in [Saxenhuber_Ramlau_2016, Saxenhuber_2016], here applied with a separated tip-tilt reconstruction in the mixed case.
In the community of AO, it is common to validate the reconstruction quality using the so called Strehl ratio into certain directions [Roddier1999]. In our simulations, we use $25$ directions positioned in a $5\times 5$ grid over the FoV. The Strehl ratio is defined as the ratio between the maximum of the real energy distribution of incoming light in the image plane $I(x,y)$ over the hypothetical distribution $I_{D}(x,y)$, which stems from the assumption of diffraction-limited imaging, i.e.,
$\text{SR}:=\frac{\max_{(x,y)}I(x,y)}{\max_{(x,y)}I_{D}(x,y)}\,.$
@@ 5.2 Numerical Results
As a first step of performance evaluation, we consider not a whole end-to-end AO simulation, but compare only the tomographic reconstruction of the proposed methods. We simulate a $3$ layer atmosphere (Figure 5.3, top row) and perform its reconstruction with the SVTD (Figure 5.3, middle row) and the iterative FD (Figure 5.3, bottom row) methods. We observe that both algorithms provide a very good result at the lowest layer $\ell=1$, but fail to reconstruct details at higher altitudes, in particular around the borders. This behaviour is expected, and generally observed for atmospheric tomography, since the reconstruction quality depends on the number of overlapping regions, which decreases with height and with a larger distance from the center, which is often referred to as field off-axis position [Ramlau_Stadler_2024].
@@
Since all algorithms yield promising tomographic reconstructions, we next consider full end-to-end AO simulations and evaluate the short-exposure (SE) and long-exposure (LE) Strehl ratio. The SE Strehl ratio provides a measure for the reconstruction quality at the current time step, whereas the LE Strehl ratio is an average over time [Roddier1999]. We compare the SVTD and FD approaches with FEWHA and the Gradient method, which are both known to provide very good results for the considered test configurations. For all results with the SVTD method, the regularization parameter $\alpha$ and the smoothing parameter $s$ were tuned experimentally a-priori. The results for the iterative FD and the Gradient method use $5$ iterations, whereas FEWHA uses $4$ iterations in its internal CG method. The method specific parameters for FEWHA have been tuned via simulations.